Circular No. 5693

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1993A AND 1993B IN ANONYMOUS GALAXIES J. Maza, University of Chile (UC); and M. Hamuy, Cerro Tololo Inter-

american Observatory, report the discovery of two supernovae on 20-min IIa-O exposures taken by G. Valladares with the CTIO Curtis Schmidt telescope. SN 1993A, at m_{pg} ∼ 18.5 on Jan. 17.17 UT, was discovered by M. Wischnjewsky (UC) and is located 5′′.5 west and 3′′.5 south of the nucleus of a galaxy at α = 7^{h}38^{m}38^{s}.4, δ = −61^{o}56′11′′ (equinox 1950.0); the supernova was not visible (m_{pg} > 20) on a IIIa-J plate taken on 1992 Dec. 29. Confirmation of SN 1993A was made by C. Anguita (UC) with a CCD V image obtained on Jan. 20.24 with the CTIO 1.5-m telescope.

SN 1993B, at m_{pg} ∼ 18 on Jan. 17.22, was discovered by R. Antezana (UC) and is located ∼ 1′′ east and 5′′ north of the nucleus of a spiral galaxy at α = 10^{h}32^{m}35^{s}.1, δ = −34^{o}11′03′′; SN 1993B was not visible (m_{pg} > 19) on a IIa-O plate taken on 1992 Dec. 27. Anguita also confirmed SN 1993B on CCD B and V frames obtained on Jan. 20.30.

COMET HELIN-LAWRENCE (1992q)

Ephemeris extension from orbital elements on MPC 21235:

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 13 23^{h}08.^{m}11 −53^{o}25.^{′}3 2.646 2.162 50.^{o}7 20.^{o}6 14.0
23 23 14.25 −53 49.8 2.662 2.126 47.6 20.0 13.9
Feb. 2 23 23.17 −54 25.2 2.652 2.095 46.3 19.9 13.8
12 23 34.86 −55 14.6 2.617 2.071 46.8 20.3 13.8
22 23 49.58 −56 20.9 2.559 2.054 49.2 21.4 13.7
Mar. 4 0 07.92 −57 47.2 2.479 2.043 53.0 22.8 13.6
14 0 30.96 −59 35.3 2.383 2.038 58.0 24.4 13.5
24 1 00.63 −61 45.1 2.275 2.041 63.8 26.0 13.4
Apr. 3 1 40.04 −64 11.1 2.163 2.051 70.1 27.3 13.3
13 2 33.84 −66 34.9 2.054 2.067 76.6 28.2 13.2
23 3 46.59 −68 15.9 1.961 2.089 82.9 28.5 13.2
May 3 5 15.44 −68 10.9 1.893 2.118 88.3 28.4 13.1
13 6 43.96 −65 34.2 1.861 2.153 92.3 28.0 13.2
23 7 56.30 −60 45.4 1.874 2.193 94.1 27.4 13.3
June 2 8 50.22 −54 47.8 1.934 2.239 93.6 26.9 13.4

1993 January 21 Daniel W. E. Green